# ORBIT: Oak Ridge Base Foundation Model for Earth System Predictability

### v1

@@ V-E Pre-Training Loss
Pre-training loss for all four model sizes is presented in Fig. 8. All four models used a fixed global data batch size of 2,880 and trained for 2.5 epochs. Despite of high initial loss, the 10 billion and 113 billion parameter models converged rapidly and outperformed the smaller 100 million and 1 billion models after seeing 2 million observation data points.
V-F Model Prediction
Figure 9: Latitude-Weighted Anomaly Correlation Coefficient (wACC) and Latitude-Weighted Root Mean Squared Error (wRMSE) scores for prediction results in the test year 2020, comparing the ORBIT models with 100 million and 1 billion parameters across four output variables.
Fig. 9 presents a comparative analysis of prediction performance, measured by wACC and wRMSE, between fine-tuned models with different parameter scales - 100 million and 1 billion. The upper panels show the model performance for z500 and t850, while the lower panels illustrate outcomes for t2m and u10. The results reveal that the 1 billion parameters models demonstrate superior prediction skills by consistently producing higher wACC and lower wRMSE across all evaluated lead times and meteorological variables. The bar graphs, depicting the wRMSE show a reduction in prediction errors for the 1 billion parameters models as opposed to the 100 million parameters models, demonstrating the advantage of employing models with increased capacity. Moreover, the line plots for wACC highlight that despite a general decline in wACC over extended forecast periods, particularly for measures such as z500 and u10, the 1 billion parameters model exhibits great stability in predictive performance, even for longer lead times up to 10 days. The sustained level of performance of the 1 billion parameters model over 100 million parameters model suggests that larger parameter scale models possess more robust and reliable forecasting capability for medium to long-range predictions and display greater potential in advancing Earth system predictability.
In this study, our main objective was to address the scaling challenges inherent in the pre-training ViT foundation model for Earth system predictability, a critical bottleneck that has not been sufficiently tackled in previous research. Our ORBIT model’s design specifically targets these scaling issues; preliminary prediction accuracy results after a few epochs already demonstrate promising results with large model sizes. This underscores the potential of our approach in effectively managing large-scale model pre-training. In the current phase of our research, we have concentrated on evaluating models with up to 1 billion parameters to establish a robust baseline for prediction accuracy across various scales and meteorological conditions. The decision to not yet compare the 10 billion and 113 billion parameter models is primarily due to time constraints and the progressive nature of our scaling experiments. This stepwise approach ensures that our findings are both scalable and reproducible within the limits of our allocation, setting a solid foundation for future expansions in model complexity. It’s important to note that models such as ClimaX, Pangu-Weather, and FourCastNet leverage high-resolution data and longer sequence processing, presenting additional complexities in direct comparison due to the variations in data scale and intricacy. In our future work, we will homogenize the data and experimental settings to compare ORBIT with the smaller parameter scale weather forecasting models in a rigorous evaluation, ensuring a fair assessment of its performance relative to existing solutions. This strategy allows us to focus initially on overcoming scalability challenges and the scaling advancements necessary for pre-training large ViTs.
@@ VI Implications
With the ORBIT model’s capacity to use large ViT architecture, incorporate 91 variables/channels, and pre-train on 10 model datasets from the CMIP6 project, a substantial advancement in AI-based Earth system modeling has been achieved. The model’s ability to process large ViT model and a wide variety of input channels allows it to capture more complex interactions and feedback mechanisms within the Earth’s climate system. This enhanced model complexity is crucial for improving the accuracy of predictions related to weather forecasting, sub-seasonal to seasonal predictions, and long-term climate projections. As climate change continues to affect global ecosystems and human societies, precise predictions become increasingly vital for planning and adaptation strategies.
The ORBIT model’s effective scaling up to 113 billion parameters on a non-NVIDIA platform, such as the Frontier supercomputer, also marks a considerable advance in the use of HPC resources for complex AI tasks. With up to 113 billion parameters, ORBIT represents the largest dense vision transformer to date, surpassing the existing largest dense vision transformer by five-fold [7], and a thousandfold larger than the largest foundational climate model [6]. This success demonstrates that the innovation in scaling algorithms can overcome challenges such as less advanced software stacks and limited interconnect capabilities. It also promotes a more inclusive approach to HPC system development that does not depend exclusively on one type of hardware/software stack, thereby enabling diversity in hardware utilization and encouraging the development of more adaptable and robust HPC systems.
Furthermore, the introduction of Hybrid Sharded Tensor-Data Orthogonal Parallelism in ORBIT showcases the deeper integration of AI and HPC. This approach optimizes the distribution of model parameters across multiple GPUs, boosting the efficiency and scalability of AI models. Such progress is essential for the future of AI applications in science and industry, where the demand for processing large-scale, complex datasets continues to grow. Efficiently scaling AI models on diverse HPC platforms will likely drive further innovations and a broader adoption of AI technologies across various domains. Importantly, this approach is not reliant on specialized HPC architectures, making it applicable on a broad range of HPC systems.
@@ Acknowledgments
The authors thank Verónica G. Melesse Vergara and Mallikarjun (Arjun) Shankar for their support of high performance computing resources. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory (ORNL), which is supported by the Office of Science of the U.S. Department of Energy (DOE) under Contract No. DE-AC05-00OR22725. This research was supported by the AI Initiative Project sponsored by the Director’s Research and Development Program in ORNL and by the BER-ASCR SciDAC Program in the DOE. This manuscript has been co-authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.

### v3

@@ V-E Pre-Training Loss
Pre-training loss for all four model sizes is presented in Fig. 8. All four models used a fixed global data batch size of 2,880 and trained for 2.5 epochs. Despite of high initial loss, the models with 10 billion and 113 billion parameters demonstrated greater training data efficiency and converged more rapidly, outperforming the smaller models with 115 million and 1 billion parameters after processing 2 million observation data points.
V-F Fine-Tuning Model Prediction
Figure 9: Latitude-Weighted Anomaly Correlation Coefficient (wACC) scores for ORBIT (115 million parameters), ClimaX (same model size), Stormer, FourCastNet, and IFS models in forecasting four atmospheric variables at 1, 14, and 30-day intervals for the testing year 2020.
Fig. 9 illustrates the performance of the 115 million-parameter ORBIT model in forecasting four key atmospheric variables at 1, 14, and 30-day intervals using the ERA5 fine-tuning dataset. This performance is compared with ClimaX [7], Stormer [4], FourCastNet [1], and IFS [28] using the wACC accuracy score. ClimaX is an AI foundation model providing forecasts up to 30 days, and its wACC performance scores were obtained from Fig. 6 of the ClimaX paper [7]. In contrast, Stormer and FourCastNet are not foundation models, but are task-specific AI models trained on ERA5 data with 1.40625 degree and 0.25 degree spatial resolutions, respectively. Their wACC performance scores were obtained from Fig. 9 of the Stormer paper [4] and Fig. 6 of the FourCastNet paper [1]. IFS is a numerical model from the European Center for Medium-Range Weather Forecasting, with its wACC performance scores obtained from Fig. 6 of the ClimaX paper [7]. Notably, Stormer provides forecasts up to 14 days, while FourCastNet and IFS do not offer either 14 or 30-day forecasts.
Fig. 9 highlights the strength of ORBIT, especially evident at longer lead times, where it consistently excels and outperforms ClimaX, Stormer, and IFS. For a 14-day forecast, ORBIT shows improvements of up to 52% over IFS and 166% over Stormer. For the 30-day forecast, ORBIT performs up to 9% better than ClimaX. This superior performance at extended lead times highlights ORBIT’s robustness and reliability in long-term forecasting. For 1-day forecasts, ORBIT’s accuracy is comparable to that of IFS, Stormer, and FourCastNet, and Climax on almost all variables except u10. ORBIT shows a 0.1% better wACC on t850 prediction compared to Climax, Stormer, and FourCastNet. It also shows a 0.27%-1.30% improvement in t2m predictions compared to all others, except IFS. For u10 forecast, ORBIT achieves a 1.26% better wACC than Climax.
It is important to note that both FourCastNet and Stormer were trained and fine-tuned on ERA5 data, with FourCastNet using a higher spatial resolution of 0.25 degree. ClimaX was fine-tuned on each atmospheric variable separately with specifically tailored hyperparameters. In contrast, ORBIT was pre-trained on CMIP6 model simulation data and fine-tuned on ERA5 data at a lower resolution of 1.40625 degree, predicting all four atmospheric variables together as a single task. Despite these differences, ORBIT still demonstrates competitive performance in 1-day forecasts and superior performance in longer-time forecasting compared to other state-of-the-art models.
V-G Fine-Tuning Data Efficiency
Figure 10: The number of the ERA5 samples processed by ORBIT fine-tuning models with different sizes until the test score converged for a 30-day prediction. As the model size increases, the number of samples needed to process is decreasing, indicating improved data efficiency with larger models.
In Fig. 10, we investigate fine-tuning data efficiency in relation to model sizes. We count the number of samples processed by the ORBIT models of three different sizes (113 million, 1 billion, and 10 billion parameters) until the test score converged to similar values. We use wACC as the test measurement and count the number of the ERA samples needed to converge during 30-day fine-tuning task. Our results demonstrate that as model size increases, the number of samples needed decreases, indicating greater fine-tuning data efficiency with larger models. The 115 million parameter model required about 76,000 samples to converge, while the 1 billion parameter model needed around 47,000 samples. Remarkably, the 10 billion parameter model achieved convergence with just 32,800 samples. This represents a significant reduction in the number of samples required, by approximately 38% for the 1 billion parameter model and 57% for the 10 billion parameter model compared to the smallest 115 million parameter model. These findings highlight a critical advantage of using larger models in terms of fine-tuning data efficiency. The reduction in the number of samples required for convergence as model size increases suggests that larger models are more effective in learning from smaller amounts of data. This efficiency can potentially lead to significant cost and time savings in fine-tuning processes, making it a valuable consideration for deploying large-scale AI models in various down stream climate and weather applications.
@@ VI Implications
Notably, the ORBIT’s exaFLOP computing throughput performance marks a significant achievement for AI-based Earth system modeling. This capability provides the foundation for a rapid computing framework that supports ORBIT’s large model size, extensive number of channels, and large number of training datasets. With the ORBIT model’s capacity to use large ViT architecture with up to 113 billion parameters, incorporate 91 variables/channels, and pre-train on 10 model datasets from the CMIP6 project, a substantial advancement in AI-based Earth system modeling has been achieved. The model’s ability to process large ViT model and a wide variety of input channels allows it to capture more complex interactions and feedback mechanisms within the Earth’s climate system. This enhanced model complexity is crucial for improving the accuracy of predictions related to weather forecasting, sub-seasonal to seasonal predictions, and long-term climate projections. As climate change continues to affect global ecosystems and human societies, precise predictions become increasingly vital for planning and adaptation strategies.
The ORBIT model’s effective scaling up to 113 billion parameters on a non-NVIDIA platform, such as the Frontier supercomputer, and with up to 1.6 exaFLOP sustained computing throughput also marks a considerable advance in the use of HPC resources for complex AI tasks. With up to 113 billion parameters, ORBIT represents the largest dense vision transformer to date, surpassing the existing largest dense vision transformer by five-fold [8], and a thousandfold larger than the largest foundational climate model, ClimaX [7], and two hundred times larger than domain-specific Stormer model [4]. This success demonstrates that the innovation in scaling algorithms can overcome challenges such as less advanced software stacks and limited interconnect capabilities. It also promotes a more inclusive approach to HPC system development that does not depend exclusively on one type of hardware/software stack, thereby enabling diversity in hardware utilization and encouraging the development of more adaptable and robust HPC systems.
Furthermore, the introduction of Hybrid Sharded Tensor-Data Orthogonal Parallelism in ORBIT showcases the deeper integration of AI and HPC. This approach optimizes the distribution of model parameters across multiple GPUs, boosting the efficiency and scalability of AI models and achieving 1.6 exaFLOPs sustained computing throughput. Such progress is essential for the future of AI applications in science and industry, where the demand for processing large-scale, complex datasets continues to grow. Efficiently scaling AI models on diverse HPC platforms will likely drive further innovations and a broader adoption of AI technologies across various domains. Importantly, this approach is not reliant on specialized HPC architectures, making it applicable on a broad range of HPC systems.
@@ Acknowledgments
The authors thank Verónica G. Melesse Vergara, Mallikarjun (Arjun) Shankar and Bronson Messer for their support of high performance computing resources. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory (ORNL), which is supported by the Office of Science of the U.S. Department of Energy (DOE). This research was primary supported by the AI Initiative Project sponsored by the Director’s Research and Development Program at ORNL, additionally supported by the BER-ASCR SciDAC Program in the DOE, and by Dan Lu’s DOE Early Career Project sponsored by the DOE BER program. This manuscript has been authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.